MEDICATION CALCULATION #73

Order: naloxone (Narcan) IV, STAT. Child's weight: 20 kg. Child's drug dosage: 0.01 mg/kg. Available: Naloxone HCI injection 400 mcg/ml. (0.4 mg/ml) 10 multidose vial.

a. How many mg should the child receive?

b. How many mL of naloxone should be administered?

Please show your step by step calculations.
Answer #1

ANSWER

    1. Amount of drug (mg) the child should receive.

                Child's drug dosage = 0.01mg / kg.

                 Child's body weight = 20 kg.

              So the amount of drug child should receive = 0.01 * 20

                                                                                    = 0.1mg.

    2. Ml of naloxone should be administered.

       Formula

         * Drug should be administered = Dose ordered (mg) * Quantity(ml) / Have(mg).

                  Dose ordered = 0.01 mg /kg ( 0.1 mg /20kg )

                     Quantity = 1 ml.

                            Have = 0.4 mg.

              As per formula

                                         = 0.1 * 1 / 0.4

                                         = 0.1 / 0.4

                                         = 0.25 ml (0.1 mg)